Show whether the differential equation given below is linear or nonlinear.
Also determine its order:
dR/dt=-k/R^n
Where n=1
1
Expert's answer
2020-12-01T20:58:44-0500

Given "\frac{dr}{dt} = \frac{-k}{R^n}" , where n = 1, therefore we have that "\frac{}{}"

"\frac{dR}{dt} = -kR^{-1}"

Since the power of the dependent variable is -1, then the differential equation is non-linear

Also the order of the differential equation is 1 since the highest derivative is 1